gsmith2002:

The product of two numbers is 713. The first number is 9 more than the second number. Which equation can be used to find x, the greater number? x(x − 9) = 713 x(x + 9) = 713 x2 − 9 = 713 x2 + 9 = 713

tigerlover:

Let x = first number and y = second number. The first statement is expressed as x*y = 713 The second statement is expressed as x = y + 9 To find x, use the second equation in terms of y, that is y = x - 9. Then substituting it to the first equation yields x*(x - 9) = 713